# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v2

@@ D.2 Detailed Analysis in Driving Comfort
Method PDM-Lite Beibwenger2024pdmLite PlanT Renz2022PlanT
@@
Pluto cheng2024pluto 56.45 $\pm$ 4.14 -0.16 $\pm$ 3.72 50.26 $\pm$ 2.17 -0.42 $\pm$ 3.38
PPO chen2024frea 74.76 $\pm$ 2.71 -0.51 $\pm$ 4.61 74.90 $\pm$ 1.21 0.40 $\pm$ 4.83
FREA chen2024frea 72.40 $\pm$ 1.72 0.29 $\pm$ 4.61 73.48 $\pm$ 3.83 -0.15 $\pm$ 4.91
FPPO-RS chen2024frea 68.33 $\pm$ 1.90 -0.07 $\pm$ 3.96 66.67 $\pm$ 0.82 -0.15 $\pm$ 3.95
@@
RS-Pluto peng2024improving 70.31 $\pm$ 4.07 0.32 $\pm$ 4.12 65.18 $\pm$ 2.11 -0.16 $\pm$ 4.07
RTR-Pluto zhang2023rtr 55.58 $\pm$ 4.76 -0.19 $\pm$ 3.37 45.12 $\pm$ 2.66 -0.14 $\pm$ 3.34
@@
Main Results. The quantitative results of the comfort metrics are presented in Table˜7. All CBV methods exhibit notable levels of driving discomfort. Although the more conservative methods identified in Section˜5.2 achieve relatively lower discomfort levels, a high baseline of discomfort persists across methods.
To investigate the underlying causes of discomfort, we further decouple the planned trajectories from the executed control actions. In CARLA, most CBV methods rely on PID controllers to transform high-level trajectory waypoints into executable driving commands, including throttle, steering, and brake. As shown in Figure˜6, the upper panel illustrates the speed tracking curve, while the middle panel presents the raw throttle signal and corresponding acceleration profile.
Because trajectory generation is performed state-wise—predicting only the immediate next action—the reference states may vary discontinuously over time. These discontinuities are amplified by the PID controller, whose binary throttle/brake responses induce abrupt changes in acceleration, ultimately leading to discomfort during vehicle operation. Such execution-level instabilities are a major contributor to the discomfort observed across CBV methods.
While many CBV methods attempt to mitigate discomfort through fine-tuning strategies that incorporate post-action feedback via reward shaping or expert action alignment—RIFT adopts a different approach. It employs a state-wise reward model (see Section˜B.7) that quantifies comfort within the trajectory’s virtual forward simulation.
To further analyze this, we visualize both the actual acceleration after executing a selected trajectory and the corresponding virtual-time acceleration (shown in Figure˜6). The results reveal that while virtual-time acceleration aligns with actual motion at the beginning of the trajectory, it underestimates acceleration variations in later segments. This leads to an overly conservative estimation of trajectory-level discomfort, resulting in insufficient supervision during training and reflected in RIFT’s comfort performance in Table˜7.
@@ D.3 Visualization of the AV-Centric Closed-Loop Simulation
To qualitatively evaluate the robustness of RIFT across diverse AV-centric scenarios, we provide additional temporal visualizations of closed-loop simulations. As shown in Figure˜7, the traffic scene consists of the autonomous vehicle (AV, controlled by PDM-Lite), background vehicles (BVs), and critical background vehicles (CBVs), which interact dynamically over time.
@@
Appendix E Social Impact
@@
This work introduces a practical framework that bridges realism and controllability in traffic simulation. By decoupling the pre-training and fine-tuning stages, our method enables models pre-trained on real-world datasets to be effectively adapted to physics-based simulators, preserving trajectory-level realism and route-level controllability while improving long-horizon, closed-loop performance. This paradigm establishes a viable pathway for transitioning data-driven approaches to physics-based simulators, facilitating more reliable closed-loop testing and training. Consequently, it contributes to the development of safer and more robust autonomous systems.

### v3

@@ D.2 Detailed Analysis in Driving Comfort
Method PDM-Lite PlanT
@@
Pluto 56.45 $\pm$ 4.14 -0.16 $\pm$ 3.72 50.26 $\pm$ 2.17 -0.42 $\pm$ 3.38
PPO 74.76 $\pm$ 2.71 -0.51 $\pm$ 4.61 74.90 $\pm$ 1.21 0.40 $\pm$ 4.83
FREA 72.40 $\pm$ 1.72 0.29 $\pm$ 4.61 73.48 $\pm$ 3.83 -0.15 $\pm$ 4.91
FPPO-RS 68.33 $\pm$ 1.90 -0.07 $\pm$ 3.96 66.67 $\pm$ 0.82 -0.15 $\pm$ 3.95
@@
RS-Pluto 70.31 $\pm$ 4.07 0.32 $\pm$ 4.12 65.18 $\pm$ 2.11 -0.16 $\pm$ 4.07
RTR-Pluto 55.58 $\pm$ 4.76 -0.19 $\pm$ 3.37 45.12 $\pm$ 2.66 -0.14 $\pm$ 3.34
@@
Main Results. The quantitative results of the comfort metrics are presented in Table˜7. All CBV methods exhibit notable levels of driving discomfort. Although the more conservative methods identified in Section˜5.2 achieve relatively lower levels of discomfort, a high baseline of discomfort persists between methods.
To investigate the underlying causes of discomfort, we further decouple the planned trajectories from the executed control actions. In CARLA, most CBV methods rely on PID controllers to transform high-level trajectory waypoints into executable driving commands, including throttle, steering, and brake. As shown in Figure˜7, the upper panel illustrates the speed tracking curve, while the middle panel presents the raw throttle signal and corresponding acceleration profile.
Because trajectory generation is performed state-wise, predicting only the immediate next action, the reference states may vary discontinuously over time. These discontinuities are amplified by the PID controller, whose binary throttle/brake responses induce abrupt changes in acceleration, ultimately leading to discomfort during vehicle operation. Such execution-level instabilities are a major contributor to the discomfort observed across CBV methods.
While many CBV methods attempt to mitigate discomfort through fine-tuning strategies that incorporate post-action feedback via reward shaping or expert action alignment, RIFT adopts a different approach. It employs a state-wise reward model (see Section˜B.7) that quantifies comfort within the trajectory’s virtual forward simulation.
To further analyze this, we visualize both the actual acceleration after executing a selected trajectory and the corresponding virtual-time acceleration (shown in Figure˜7). The results reveal that while virtual-time acceleration aligns with actual motion at the beginning of the trajectory, it underestimates acceleration variations in later segments of the trajectory. This leads to an overly conservative estimation of trajectory-level discomfort, resulting in insufficient supervision during training and reflected in RIFT’s comfort performance in Table˜7.
@@ D.3 Visualization of the AV-Centric Closed-Loop Simulation
To qualitatively evaluate the robustness of RIFT across diverse AV-centric scenarios, we provide additional temporal visualizations of closed-loop simulations. As shown in Figure˜8, the traffic scene consists of the autonomous vehicle (AV, controlled by PDM-Lite), background vehicles (BVs), and critical background vehicles (CBVs), which interact dynamically over time.
@@
Appendix E Discussion and Broader Implications
E.1 Use of Large Language Models (LLMs)
The large language model (LLM) was employed as a general-purpose writing assistant during the preparation of this manuscript. Its use was limited to:
  • Language refinement: improving grammar, syntax, and overall readability to ensure clarity and professionalism.
  • Style adjustments: suggesting more concise and precise phrasing while preserving the original meaning and technical content.
The LLM was not involved in research ideation, experimental design, data collection, analysis, or interpretation of results. All intellectual contributions and scientific conclusions are solely those of the authors. This disclosure is provided in accordance with the conference guidelines on LLM usage.
E.2 Limitations and Future Work.
In the current framework, the generation head is frozen during RL fine-tuning, and its reliability stems from the robust trajectory generation capability learned during IL pre-training. However, without reliable expert demonstrations, the realism and robustness of generated trajectories cannot be further improved during RL fine-tuning. This limitation highlights a key avenue for future research: developing methods—potentially leveraging RL or other self-improvement paradigms—that can enhance the trajectory generation quality without relying on expert demonstrations.
E.3 Social Impact
@@
This work presents a practical framework that bridges the gap between realism and controllability in traffic simulation. By decoupling pre-training and fine-tuning, our method enables models pre-trained on real-world datasets to adapt effectively to physics-based simulators, preserving trajectory-level realism and route-level controllability while improving long-horizon closed-loop performance. This paradigm establishes a viable pathway for transitioning data-driven approaches to physics-based simulators, enabling more reliable closed-loop testing and training. Consequently, it advances safer and more robust autonomous systems.